Solve the system by addition. x + y = 4
x – y = 10

Answer by funmath(2933)

Solve the system by addition.
Add the two equations and the y's are eliminated. Now you can solve for x.
x + y = 4
x – y = 10
_____________
2x+0=14
2x=14
2x/2=14/2
x=7
:
Let x=7 in either equation and solve for y. I'm using the first one.
7+y=4
-7+7+y=4-7
y=-3
:
The solution is (x,y)=(7,-3)
HGappy Calculating!!!